OpenStudy (anonymous):

A curve is such athat {dy/dx}= {12 over (2x+1)^2} and P(1,5) is a point on the curve. (i) The normal to the curve at P crosses the x-axis at Q. Find the coordiantes of Q.

OpenStudy (callisto):

slope of normal = -1/ (dy/dx) you've got the slope and point so find he equation for the normal and you can find the coordinates of Q

OpenStudy (callisto):

Hmm.. normal is perpendicular to tangent. dy/dx is the slope of tangent. Do you understand what i've written?

OpenStudy (anonymous):

the slope of the normal meeting the curve = -1 so m1m2= -1 -1/dxdy= m2 (the slope of the normal) Does this equal to (2x+1)^2/12?

OpenStudy (callisto):

you've forgotten the negative sign..

OpenStudy (anonymous):

Oh, yes. - (2x+1)^2 /12

OpenStudy (anonymous):

Then I have to find the equation of the line? How?

OpenStudy (callisto):

point - slope form

OpenStudy (anonymous):

y=0 no?

OpenStudy (callisto):

(y-y1) = m(x-x1) (x1,y1) = (1,5)

OpenStudy (callisto):

m = - (2x+1)^2 /12 and put x=1into m=...

OpenStudy (anonymous):

Hm, Ok... m= -(2x+1)^2/12 m= -9/12 m= -3/4?

OpenStudy (callisto):

should be like that

OpenStudy (anonymous):

So, then y=-3/4x+b Then do I plug in the y and x values of (1,5)?

OpenStudy (callisto):

(y-y1) = m(x-x1) (x1,y1) = (1,5) Sorry if i've used the wrong words...

OpenStudy (anonymous):

I mean, would the equation of the line be y= -3/4x+5 3/4?

OpenStudy (callisto):

shouldn't it be y= -3/4x+4 3/4?

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

the m value is -3/4, so if you move that, it's 5+3/4

OpenStudy (callisto):

Sorry my mistakes :(

OpenStudy (anonymous):

That's OK... So then do you plug in 0 in the y value?

OpenStudy (anonymous):

to get x, when it's Q

OpenStudy (callisto):

yup!

OpenStudy (anonymous):

Can you see what you get and so I can double check my answer?

OpenStudy (anonymous):

I got 7.67 = x?

OpenStudy (callisto):

something ugly, 7 2/3, 7.67 if you need a decimal number

OpenStudy (anonymous):

Yes. It looks ugly... I hope it's correct! :)

OpenStudy (anonymous):

The next question says find the equation of the curve?....?

OpenStudy (callisto):

use integration

OpenStudy (callisto):

\[y=\int\limits dy/dx\]

OpenStudy (anonymous):

I thought we have the slope of the curve, which is 12/(2x+1)^2? So, can't we just use that? Why do we have to integrate?

OpenStudy (callisto):

Ehhh.. you need the curve right?

OpenStudy (anonymous):

I need the equation of the curve...

OpenStudy (callisto):

then you need to know that dy/dx is the slope of the curve, to find the curve you need to do the reverse..

OpenStudy (anonymous):

Oh... I just thought you could use that slope and do y=mx+b, where m is the slope??

OpenStudy (callisto):

nope... it's a curve, not a straight line. For a curve, the slope differs from point to point. So, there are no constant slope

OpenStudy (callisto):

*there is

OpenStudy (anonymous):

Oh, Ok. So.... how do you integrate it?

OpenStudy (callisto):

I'm not good at using math latex here.. give me some time to write it..

OpenStudy (anonymous):

Sure. I'll be back in a moment.

OpenStudy (anonymous):

am back

OpenStudy (anonymous):

Thank you! So, is that the equation?

OpenStudy (anonymous):

Wait, why is it timesed by 1/2?

OpenStudy (callisto):

Personally, i think it is, but not sure

OpenStudy (anonymous):

Should I ask Sam?

OpenStudy (callisto):

you can d/dx(equation) and see if you understand why it is multiplied by 1/2

OpenStudy (callisto):

Sure,you should , I though he should have taught you since you asked the question.. but that falls short of my expectation :(

OpenStudy (anonymous):

@.Sam. Can you check all these answers.?

OpenStudy (anonymous):

Callisto, there's a last part of the question.

OpenStudy (callisto):

Honestly, i've learnt differentiation for a year and integration for half a year.. So , i'm not really good at that :S (no time to practise)

OpenStudy (anonymous):

Um... the last part of the question says, A point is moving along the curve in such a way that the x=coordinate is increasing at a constant rate of 0.3 units per second. Find the rate of increase of the y-coordinate when x=1

OpenStudy (anonymous):

your integretion was correct :)

OpenStudy (anonymous):

Um, I have to go very soon. Can you see if you can help me with the last question, and then it'll be good?

OpenStudy (callisto):

Sorry was away from computer. I'm not sure for this problem from the question, you'll have dx/dt = 0.03 and you can do it like that dy/dt = (dy/dx)(dx/dt) to find dy/dt..

OpenStudy (callisto):

sorry again, it should be dx/dt = 0.3

OpenStudy (callisto):

then it's like dy/dt = {12 over (2x+1)^2} * 0.3 Put x=1into dy/dt and you can find the answer, i think